OpenStudy (anonymous):

determine the real roots : (5x^2+20) (3x^2-48)

OpenStudy (anonymous):

since the numerator is what will cause y=0 you want the numerator to = 0 \[5x^2+20=0\] \[5x^2=-20\] \[x^2=-4\] \[x=\sqrt{-4}\] \[x=\pm 2i\]

OpenStudy (anonymous):

there is only complex roots

OpenStudy (anonymous):

thaank you (:

OpenStudy (anonymous):

I don't think that's the question.

OpenStudy (anonymous):

the answer is + 4 or -4

OpenStudy (anonymous):

y =(5x^2+20)(3x^2-48) has real roots at x = +/- 4

OpenStudy (anonymous):

how did you solve for that? O:

OpenStudy (anonymous):

roots when y = 0 ie when (x^2+4) = 0 or (x^2 - 16) = 0

OpenStudy (anonymous):

oh i saw it as division... -.-

OpenStudy (anonymous):

yes the second part would be \[3(x^2-16)\]

OpenStudy (anonymous):

\[=(x+4)(x-4)\]

OpenStudy (anonymous):

x^2 = -4 x= sqrt(-2) x= +/- 2i *imaginary root* x^2 =16 x =sqrt(4) = +/- 4

OpenStudy (anonymous):

\[(5x^2+20)(3x^2-48)\] \[5(x^2+4)3(x^2-16)\] \[15(x^2+4)(x^2-16)\] \[15 \neq 0\] \[x^2+4=0\] \[x^2=-4\] \[x=\pm 2i\] \[x^2-16=0\] \[x^2=16\] \[x=\pm 4\] \[15(x-2i)(x+2i)(x-4)(x+4)\]

OpenStudy (anonymous):

what does the i stand for ?

OpenStudy (anonymous):

nvm i get it (: thank you